# SFormer-OR: Single-Stage Bimodal Transformer for Scene Graph Generation in OR

### v1

@@ Abstract
Scene graph generation (SGG) of surgical procedures is crucial in enhancing holistically cognitive intelligence in the operating room (OR). However, previous works have primarily relied on the multi-stage learning that generates semantic scene graphs dependent on intermediate processes with pose estimation and object detection, which may compromise model efficiency and efficacy, also impose extra annotation burden. In this study, we introduce a novel single-stage bimodal transformer framework for SGG in the OR, termed SFormer-OR, aimed to complementally leverage multi-view 2D scenes and 3D point clouds for SGG in an end-to-end manner. Concretely, our model embraces a View-Sync Transfusion scheme to encourage multi-view visual information interaction. Concurrently, a Geometry-Visual Cohesion operation is designed to integrate the synergic 2D semantic features into 3D point cloud features. Moreover, based on the augmented feature, we propose a novel relation-sensitive transformer decoder that embeds dynamic entity-pair queries and relational trait priors, which enables the direct prediction of entity-pair relations for graph generation without intermediate steps. Extensive experiments have validated the superior SGG performance and lower computational cost of SFormer-OR on 4D-OR benchmark, compared with current OR-SGG methods, e.g., 3% Precision increase and 24.2M reduction in model parameters. We further compared our method with generic single-stage SGG methods with broader metrics for a comprehensive evaluation, with consistently better performance achieved. The code will be made available.
Scene graph generation, 3D surgical scene understanding, Transformer, single-stage, bimodal.
@@ 1 Introduction
[…]
To advance SGG methods for OR modeling, Özsoy et al. [6] introduced the first OR-SGG dataset, termed 4D-OR, which contains multi-view video frames and corresponding 3D point clouds in the knee surgery theatre. Moreover, they proposed a multi-stage semantic SGG pipeline that provides 3D human pose estimation and object detection, followed by blending different forms of prediction labels and constructing a scene graph through a 3D SGG framework [7]. Subsequently, LABRAD-OR [5] proposed to enhance the multi-stage framework by incorporating temporal information into scene graphs. The constructed memory scene graphs can capture fine-grained relationships between entities in dynamic OR, achieving state-of-the-art results. However, existing methods rely on a multi-stage model training (shown in Fig. 1 (a)), typically facing the following challenges which may hinder its applicability in real-world deployment: i) In terms of efficiency, multi-stage approaches require step-wise processing, which imposes additional annotation burdens and computational complexity during model training. Meanwhile, the inherent structure of a multi-stage framework elongates inference time, potentially compromising the ability to detect abnormalities in OR promptly. ii) In terms of precision, the intermediate representations generated at each stage may not fully retain the contextual information desired for scene modeling, consequently failing to capture the long-range dependencies within the scene and risking losing representational information. To overcome these limitations, we aim at the single-stage paradigm that properly utilizes both 2D images and 3D point cloud modalities to generate OR scene graphs.
@@
On the other hand, typical 3D SGG approaches [7, 14, 15] on point cloud data assume the availability of 3D instance labels as known inputs. This assumption significantly increases annotation efforts and is often impractical in a clinical setting. Moreover, surgical actions performed by the chief surgeon are very subtle, making it difficult for 3D SGG methods to precisely capture specific operations without sufficient semantic information. In this regard, how to efficiently leverage geometric information and productively dig out 2D semantic cues is crucial for accurate scene graph generation in the OR. Concentrating on this issue, we strive to design a concise multimodal SGG paradigm for OR scenarios.
In this paper, we present a novel single-stage multi-view bimodal transformer-based SGG model, termed SFormer-OR, for efficient scene graph generation in OR recordings. As illustrated in Fig. 1 (b), our single-stage end-to-end SGG paradigm enables the efficient generation of scene graphs, providing a more concise and effective solution for scene understanding in surgery. In contrast to multi-stage OR-SGG methods [6, 5], our model requires no integration of intermediate predicted instances and has no reliance on the supervision of pose estimation labels. Concretely, we propose a relation-sensitive transformer in the relation decoder of SFormer-OR. This structure first couples the output embeddings produced by the transformer decoder into a set of subject-object pairs and then incorporates relational trait priors to generate dynamic relation queries. By interacting with multi-view memories through cross-attention, our model predicts relationships within subject-object pairs and achieves the final scene graph. To better converge and exploit features from different modalities, a geometry-visual cohesion (GVC) module is introduced to assemble 2D multi-view semantic cues and 3D geometric features to generate appearance-structure representations, instead of the general projection operation in most existing works. In the multi-view unit, we also present a view-sync transfusion (VST) scheme that complements the local fine-level information of multiple views to the main view features to enhance the holistic visual semantic representation and mitigate occlusions. Extensive experimental findings demonstrate the effectiveness of ‘single stage’.
@@
To our best knowledge, SFormer-OR is the first single-stage bimodal framework for SGG in the operating rooms. The proposed OR-SGG model can directly generate accurate scene graphs with end-to-end training.
@@
Extensive validations demonstrate that our SFormer-OR achieves superior performance on the 4D-OR benchmark, whether compared to multi-stage OR-SGG methods or to generic single-stage SGG models.
@@ 2.1 Semantic Scene Graph in OR
The main literature in surgical data science focuses on analyzing the surgical scene, such as surgical workflow and tool recognition [16], instrument segmentation [17], and scene understanding [18]. Scene understanding in dynamic operating room (OR) environments is a new challenging task, which builds entity interactive systems by semantically modeling the holistic space [6, 5]. Özsoy et al. [6] first revealed the challenge of SSG generation for the OR scene. Specifically, they built a new scene graph dataset (4D-OR) for knee surgery scenarios, which consists of multi-view RGB frames and point cloud data accompanied by human pose labels, 3D object bounding boxes, and scene graph labels. Accordingly, they also proposed a multi-stage SGG baseline, which first predicts the human pose and object proposals by VoxelPose [19] and Group-Free [20] respectively, and then combines the prediction labels and feeds them into the 3DSSG model [7] for scene graph prediction. Building upon the 4D-OR baseline, LABRAD-OR [5] is a novel lightweight SGG model that introduces the concept of memory scene graphs. This model fuses temporal cues with visual features to improve the SGG accuracy. However, both approaches are based on multi-stage architectures, while real clinical scenarios require more on the real-time performance that these models lack, instead of their multi-stage outputs. To this end, we propose a single-stage bimodal SGG framework. As depicted in Fig. 1, our model takes multi-view inputs and point cloud data to directly generate scene graphs without intermediate processing, enhancing the practicality of OR-SGG models in real-world surgery.
@@
The main purpose of Scene Graph Generation (SGG) is to construct representations of visual relationships between objects, which has been extensively studied in the field of computer vision in 2D images [4]. In the deep learning era, mainstream SGG approaches [10, 8, 9, 11] primarily adopt a two-stage framework, firstly locating objects using an object detection network, and then employing various mechanisms to purify contextual structural information and refine location-specific features, e.g., message passing [8], attention module [11, 21], and visual translation [22], to predict relationships by classifiers and generate scene graphs. Recently, inspired by fully convolutional network (FCN) in object detection models [23], one-stage SGG methods employ FCN-based [24, 25] or CNN-Transformer architectures [12, 13] to efficiently predict visual relationships directly from global features using sparse proposal sets and queries, making this paradigm concise and easy to converge. To further improve performance, subsequent one-stage models, such as SGTR [12] and RelTR [13], utilize the Transformer architecture and the attention mechanism to capture global perception and predict triplets by interacting with subject-object queries. However, 2D images lack critical geometric information for SGG in ORs. For better capturing location-related relationships, our model incorporates 3D point cloud inputs that complement 2D semantic information from the images to enhance the perception of structural correlations.
@@
In contrast to 2D SGG at the image level, feeding point clouds to 3D SGG models provides a more robust representation of the geometric properties of objects [15], leading to a higher sensitivity in predicting structure-aware relationships. Wald et al. [7] developed a large-scale 3D semantic scene graph (SSG) dataset, termed 3DSSG. They also proposed an end-to-end SGG pipeline for constructing 3D scene graphs via graph convolution networks (GCNs). Later, Zhang et al. [14] introduced an approach for directly generating scene graphs from 3D point clouds, which encodes the point clouds into more discriminative latent spaces as a priori knowledge. Nevertheless, it is challenging to learn semantic information from the point cloud, since the point cloud is defined on geometric structures instead of more semantic colors or textures. For this reason, VL-SAT [15] incorporates 2D instance labels for multimodal training, which assists the 3D-only SGG model in capturing richer semantic information. It is worth noting that the 3DSSG benchmark generally assumes the instance labels as known inputs during model training. In contrast, for the 4D-OR benchmark, instance labeling in the operating theatre is difficult and laborious to acquire. This drives us to design a bimodal SGG model that complements 3D modalities by 2D semantics to generate a comprehensive scene understanding without the need for known instance labels.
@@ 3 Methodology
The overall architecture of the proposed single-stage bimodal SFormer-OR is described in Fig. 2. For the input of four view images, we employ a CNN-Transformer encoder to extract multi-view semantic features. Subsequently, we introduce a view-sync transfusion (VST) scheme that complements the multi-view information into the main view to attain comprehensive and enriched synergic features. Bilaterally, for 3D point cloud input, we adopt PointNet++ [26] as the point encoder to capture abundant geometric features. Given 2D and 3D modal representations, we design a geometry-visual cohesion (GVC) operation to implicitly interact synergic semantic information with point cloud features to achieve unified appearance-structure features. After that, we leverage the unified feature to yield entity (i.e., subject and object) proposals and predict relationships. In the operating room, the entities include surgeons, patients, operating tables, instruments, etc. In the entity unit, we initialize a set of entity queries and produce entity proposals with the instance decoder in line with DETR [27]. In order to predict relationships without relying on predictions from intermediate stages, we propose a relation-sensitive transformer decoder to perceive interactions between entity pairs and generate OR scene graphs. We will detail the structure of VST, GVC, and relation-sensitive transformers embedded in SFormer-OR in the following subsections.
3.1 Multi-View and Bimodal Interaction
SFormer-OR comprises two encoder units to process multi-view 2D images and 3D point cloud data of the whole scene, respectively. To utilize the multi-view representation to comprehensively emphasize surgical operations and enhance structure perception through holistic point cloud samples, we introduce a view-sync transfusion (VST) scheme to synergize multi-view features and a geometry-visual cohesion (GVC) operation to complement the geometric information.
@@ 3.2.1 Entity Predictions
Before predicting relationship triplets [entity, relation, entity] which contain the relation between subject entity and object entity in the operating theatre, we first generate entity boxes in the entity unit, allowing for the subsequent pairing of entities to predict the relationships in the surgical scene. Following DETR [27], we utilize a set of randomly initialized queries as input to the transformer decoder. In the transformer decoder, we interact the unified feature with queries to produce the output embedding via cross-attention operations. Then an FFN is used to get entity proposals with class labels. During the training phase, we supervise the prediction of entity boxes by 2D box labels. It is worth noting that box supervision is trained together with the final scene graph labels without additional pre-training.
@@ 3.2.2 Dynamic Relation Queries
The previous multi-stage SGG-OR models [6, 5] first predict different types of instance labels (i.e., human pose estimation labels and object detection labels), and then merge them through intermediate steps before feeding them into the subsequent SGG network. In contrast, our framework proposes to establish entity connections after generating entity proposals and then predicts potential relationships between entity pairs. Inspired by the concept of building edges between nodes in graph convolutional networks (GCNs), in relation unit, we construct all potential subject-object pairs. For the subject and object in each pair, we extract semantic features, spatial features, and supplemental point cloud features within the entity regions to generate relational trait priors and then yield dynamic relation queries (see Fig. 3). Specifically, we first filter out entities with class scores greater than 0.5 and pair them to generate entity pairs. Here the number of is not fixed and is dynamically changed according to the perception of entity proposals in the entity unit. To obtain corresponding semantic features, we concatenate the entity-aware embeddings associated with entities. Spatial features are calculated based on the positional attributes between entities, including differences in centroid coordinates, distances between the coordinates of center points, and bounding box areas. Given a pair of entities with center point coordinates and , the spatial feature can be computed as

### v2

# S2Former-OR: Single-Stage Bi-Modal Transformer for Scene Graph Generation in OR

@@ Abstract
Scene graph generation (SGG) of surgical procedures is crucial in enhancing holistically cognitive intelligence in the operating room (OR). However, previous works have primarily relied on multi-stage learning, where the generated semantic scene graphs depend on intermediate processes with pose estimation and object detection. This pipeline may potentially compromise the flexibility of learning multimodal representations, consequently constraining the overall effectiveness. In this study, we introduce a novel single-stage bi-modal transformer framework for SGG in the OR, termed S2Former-OR, aimed to complementally leverage multi-view 2D scenes and 3D point clouds for SGG in an end-to-end manner. Concretely, our model embraces a View-Sync Transfusion scheme to encourage multi-view visual information interaction. Concurrently, a Geometry-Visual Cohesion operation is designed to integrate the synergic 2D semantic features into 3D point cloud features. Moreover, based on the augmented feature, we propose a novel relation-sensitive transformer decoder that embeds dynamic entity-pair queries and relational trait priors, which enables the direct prediction of entity-pair relations for graph generation without intermediate steps. Extensive experiments have validated the superior SGG performance and lower computational cost of S2Former-OR on 4D-OR benchmark, compared with current OR-SGG methods, e.g., 3 percentage points Precision increase and 24.2M reduction in model parameters. We further compared our method with generic single-stage SGG methods with broader metrics for a comprehensive evaluation, with consistently better performance achieved. Our source code can be made available at: https://github.com/PJLallen/S2Former-OR.
Scene graph generation, 3D surgical scene understanding, Transformer, single-stage, bi-modal.
@@ 1 Introduction
As medicine and information technology advance, the operating room (OR) has evolved into a highly complicated and technologically rich environment. This setting is characterized by high variability and dynamic, unpredictable scenarios [1]. Surgical procedures in the modern OR are increasingly complex, necessitating intricate coordination among medical personnel, patients, and a multitude of devices [2]. An automatic and holistic comprehension of surgical procedures is critically essential for monitoring and optimizing surgical processes, scheduling surgeons, and enhancing coordination among surgical teams [3]. To this end, holistically understanding OR can augment OR efficiency, improve patient safety, and enhance the quality of patient treatment [4]. Scene graph generation (SGG), a powerful semantic representation technique derived from the computer vision field, has shown great potential in capturing complex interactive relationships between entities in the 3D dynamic OR environment [4, 5]. Effective SGG facilitates a comprehensive understanding of the entire OR scene, thus highly desirable for promoting cognitive assistance for the next generation of operating intelligence.
To advance SGG methods for OR modeling, Özsoy et al. [6] introduced the first OR-SGG dataset, termed 4D-OR, which contains multi-view video frames and corresponding 3D point clouds in the knee surgery theatre. Moreover, they proposed a multi-stage semantic SGG pipeline that provides 3D human pose estimation and object detection, followed by blending different forms of prediction labels and constructing a scene graph through a 3D SGG framework [7]. Subsequently, LABRAD-OR [5] proposed to enhance the multi-stage framework by incorporating temporal information into scene graphs. The constructed memory scene graphs can capture fine-grained relationships between entities in dynamic OR, achieving state-of-the-art results. Generally, existing OR-SGG methods based on multi-stage step-wise training (shown in Fig. 1(a)), typically confront the following challenges: i) For multimodal scene modeling, the intermediate representations produced by each stage may be inadequate in preserving the contextual information desired for scene generating, consequently being suboptimal to capturing the long-range dependencies within the scene and constraining the adaptability of representation learning. ii) The multi-stage pipeline requires step-by-step training of multiple subtasks, which may limit operational efficiency and hinder its deployment and application in realistic surgery. To overcome these limitations, we aim at the single-stage paradigm that properly utilizes both 2D images and 3D point cloud modalities to generate OR scene graphs.
@@
On the other hand, typical 3D SGG approaches [7, 14, 15] on point cloud data assume the availability of 3D instance labels as known inputs. This assumption significantly increases annotation efforts and is often impractical in a clinical setting. Moreover, surgical actions performed by the chief surgeon are very subtle, making it difficult for 3D SGG methods to precisely capture specific operations without sufficient semantic information. In this regard, efficiently leveraging geometric information and effectively extracting 2D semantic cues are crucial for accurate scene graph generation in ORs. Concentrating on this issue, we strive to design a concise multimodal SGG paradigm for OR scenarios.
In this paper, we present a novel single-stage multi-view bi-modal transformer-based SGG model, termed S2Former-OR, for efficient scene graph generation in OR recordings. As illustrated in Fig. 1(b), our single-stage end-to-end SGG paradigm enables the efficient generation of scene graphs, providing a more concise and effective solution for scene understanding in surgery. In contrast to multi-stage OR-SGG methods [6, 5], our model requires no integration of intermediate predicted instances and has no reliance on the supervision of pose estimation labels. Concretely, we propose a relation-sensitive transformer in the relation decoder of S2Former-OR. This structure first couples the output embeddings produced by the transformer decoder into a set of subject-object pairs and then incorporates relational trait priors to generate dynamic relation queries. By interacting with multi-view memories through cross-attention, our model predicts relationships within subject-object pairs and achieves the final scene graph. To better converge and exploit features from different modalities, a geometry-visual cohesion (GVC) module is introduced to assemble 2D multi-view semantic cues and 3D geometric features to generate appearance-structure representations, instead of the general projection operation in most existing works. In the multi-view unit, we also present a view-sync transfusion (VST) scheme that complements the local fine-level information of multiple views to the main view features to enhance the holistic visual semantic representation and mitigate occlusions. Extensive experimental findings demonstrate the effectiveness of ‘single stage’.
@@
To our best knowledge, S2Former-OR is the first bi-modal single-stage framework for SGG in the operating rooms. The proposed OR-SGG model can directly generate accurate scene graphs with end-to-end training.
@@
Extensive validations demonstrate that our S2Former-OR achieves superior performance on the 4D-OR benchmark, whether compared to multi-stage OR-SGG methods or to generic single-stage SGG models.
@@ 2.1 Semantic Scene Graph in OR
The main literature in surgical data science focuses on analyzing the surgical scene, such as surgical workflow and tool recognition [16], instrument segmentation [17], and scene understanding [18]. Scene understanding in dynamic operating room (OR) environments is a new challenging task, which builds entity interactive systems by semantically modeling the holistic space [6, 5]. Özsoy et al. [6] first revealed the challenge of SSG generation for the OR scene. Specifically, they built a new scene graph dataset (4D-OR) for knee surgery scenarios, which consists of multi-view RGB frames and point cloud data accompanied by human pose labels, 3D object bounding boxes, and scene graph labels. Accordingly, they also proposed a multi-stage SGG baseline, which first predicts the human pose and object proposals by VoxelPose [19] and Group-Free [20] respectively, and then combines the prediction labels and feeds them into the 3DSSG model [7] for scene graph prediction. Building upon the 4D-OR baseline, LABRAD-OR [5] is a novel lightweight SGG model that introduces the concept of memory scene graphs. This model fuses temporal cues with visual features to improve the SGG accuracy. However, both approaches are based on multi-stage architectures, while in real clinical scenarios, there is a greater demand for streamlined and easily deployable end-to-end multimodal models to simplify the modeling workflow and enhance the flexibility of applications in diverse clinical devices. To this end, we propose a single-stage bi-modal SGG framework. As depicted in Fig. 1(b), our model takes multi-view inputs and point cloud data to directly generate scene graphs without intermediate processing, enhancing the practicability of OR-SGG models in real-world surgery.
@@
The main purpose of Scene Graph Generation (SGG) is to construct representations of visual relationships between objects, which has been extensively studied in the field of computer vision in 2D images [4]. In the deep learning era, mainstream SGG approaches [10, 8, 9, 11] primarily adopt a two-stage framework, firstly locating objects using an object detection network, and then employing various mechanisms to purify contextual structural information and refine location-specific features, e.g., message passing [8], attention module [11, 21], and visual translation [22], to predict relationships by classifiers and generate scene graphs. Recently, inspired by fully convolutional network (FCN) in object detection models [23], one-stage SGG methods employ FCN-based [24, 25] or CNN-Transformer architectures [12, 13] to efficiently predict visual relationships directly from global features using sparse proposal sets and queries, making this paradigm concise and easy to converge. Subsequently, Teng et al. [26] proposed a structured sparse R-CNN framework based on set prediction to directly generate scene graphs without the need for explicit object detection and preceding graph construction. To further improve performance, subsequent one-stage models, such as SGTR [12] and RelTR [13], utilize the Transformer architecture and the attention mechanism to capture global perception and predict triplets by interacting with subject-object queries. Furthermore, Pix2SG [25] is proposed to allow the model to make multiple prediction sequences sequentially from a set of objects like an autoregressive decoder and leverages a transformer-based sequence generation design to liberate the dependence on spatial localization. However, 2D images lack critical geometric information for SGG in ORs. For better capturing location-related relationships, our model incorporates 3D point cloud inputs that complement 2D semantic information from the images to enhance the perception of structural correlations.
@@
In contrast to 2D SGG at the image level, feeding point clouds to 3D SGG models provides a more robust representation of the geometric properties of objects [15], leading to a higher sensitivity in predicting structure-aware relationships. Wald et al. [7] developed a large-scale 3D semantic scene graph (SSG) dataset, termed 3DSSG. They also proposed an end-to-end SGG pipeline for constructing 3D scene graphs via graph convolution networks (GCNs). Later, Zhang et al. [14] introduced an approach for directly generating scene graphs from 3D point clouds, which encodes the point clouds into more discriminative latent spaces as a priori knowledge. Nevertheless, it is challenging to learn semantic information from the point cloud, since the point cloud is defined on geometric structures instead of more semantic colors or textures. For this reason, VL-SAT [15] incorporates 2D instance labels for multimodal training, which assists the 3D-only SGG model in capturing richer semantic information. It is worth noting that the 3DSSG benchmark generally assumes the instance labels as known inputs during model training. In contrast, for the 4D-OR benchmark, instance labeling in the operating theatre is difficult and laborious to acquire. This drives us to design a bi-modal SGG model that complements 3D modalities by 2D semantics to generate a comprehensive scene understanding without the need for known instance labels.
@@ 3 Methodology
The overall architecture of the proposed single-stage bi-modal S2Former-OR is described in Fig. 2. For the input of four view images, we employ a CNN-Transformer encoder to extract multi-view semantic features. Subsequently, we introduce a view-sync transfusion (VST) scheme that complements the multi-view information into the main view to attain comprehensive and enriched synergic features. Bilaterally, for 3D point cloud input, we adopt PointNet++ [27] as the point encoder to capture abundant geometric features. Given 2D and 3D modal representations, we design a geometry-visual cohesion (GVC) operation to implicitly interact synergic semantic information with point cloud features to achieve unified appearance-structure features. After that, we leverage the unified feature to yield entity (i.e., subject and object) proposals and predict relationships. In the operating room, the entities include surgeons, patients, operating tables, instruments, etc. In the entity unit, we initialize a set of entity queries and produce entity proposals with the instance decoder in line with DETR [28]. In order to predict relationships without relying on predictions from intermediate stages, we propose a relation-sensitive transformer decoder to perceive interactions between entity pairs and generate OR scene graphs. We will detail the structure of VST, GVC, and relation-sensitive transformers embedded in S2Former-OR in the following subsections.
3.1 Multi-View and bi-modal Interaction
S2Former-OR comprises two encoder units to process multi-view 2D images and 3D point cloud data of the whole scene, respectively. To utilize the multi-view representation to comprehensively emphasize surgical operations and enhance structure perception through holistic point cloud samples, we introduce a view-sync transfusion (VST) scheme to synergize multi-view features and a geometry-visual cohesion (GVC) operation to complement the geometric information.
@@ 3.2.1 Entity Predictions
Before predicting relationship triplets [entitysub, relation, entityobj] which contain the relation between subject entity and object entity in the operating theatre, we first generate entity boxes in the entity unit, allowing for the subsequent pairing of entities to predict the relationships in the surgical scene. Following DETR [28], we utilize a set of randomly initialized queries as input to the transformer decoder. In the transformer decoder, we interact the unified feature with queries to produce the output embedding via cross-attention operations. Then an FFN is used to get entity proposals with class labels. During the training phase, we supervise the prediction of entity boxes by 2D box labels. It is worth noting that box supervision is trained together with the final scene graph labels without additional pre-training.
@@ 3.2.2 Dynamic Relation Queries
The previous multi-stage OR-SGG models [6, 5] first predict different types of instance labels (i.e., human pose estimation labels and object detection labels), and then merge them through intermediate steps before feeding them into the subsequent SGG network. In contrast, our framework proposes to establish entity connections after generating entity proposals and then predicts potential relationships between entity pairs. Inspired by the concept of building edges between nodes in graph convolutional networks (GCN), in relation unit, we construct all potential subject-object pairs. For the subject and object in each pair, we extract semantic features, spatial features, and supplemental point cloud features within the entity regions to generate relational trait priors and then yield dynamic relation queries (see Fig. 3). Specifically, we first filter out entities with class scores greater than 0.5 and pair them to generate entity pairs. Here the number of is not fixed and is dynamically changed according to the perception of entity proposals in the entity unit. To obtain corresponding semantic features, we concatenate the entity-aware embeddings associated with entities. Spatial features are calculated based on the positional attributes between entities, including differences in centroid coordinates, distances between the coordinates of center points, and bounding box areas. Given a pair of entities with center point coordinates and , the spatial feature can be computed as